OpenStudy (anonymous):

(a) Find the exponential growth or decay model y = ae^(bt) or y = ae^(−bt) for the population of each country by letting t = 10 correspond to 2010. Use the model to predict the population of each country in 2030. Country 2010 2020 Country A 8.7 7.2 Country B 32.9 37.3 Country C 1360.5 1422.6 Country D 62.3 65.4 Country E 305.7 343.9

OpenStudy (anonymous):

pls help:(

OpenStudy (anonymous):

Do you understand the fundamental difference between the population growth trends in either model? \[\large y=ae^{bt}\quad\text{vs}\quad y=ae^{-bt}\]

OpenStudy (anonymous):

i have no idea how to approach this problem

OpenStudy (campbell_st):

ok... so they have given you the model as well as A, y and t so to start you need to find k, the growth constant... so for country A you know you are given the initial population a = 8.7 , and then 10 years later t = 10 you are given y = 7.2 so the population is in decline.. and k will be negative. so to solve for k start with \[7.2 = 8.7 e^{10k}\] then dividing both sides by 8.7 you get \[\frac{7.2}{8.7} = e^{10k}\] to find k, you need to take the base e log of both sides \[\ln(\frac{7.2}{8.7} )= 10k\] then divide both sides of the equation to find k. \[k = \frac{\ln(\frac{7.2}{8.7})}{10}\] then when you have k, substitute it into the model along with a = 8.7 and t = 20 to find the population in 2030. hope it helps

OpenStudy (anonymous):

yes! can u help with one more question?